Part A Calculate the pH of 0.85 M NH3 (Kb=1.8×10−5). Express your answer using two decimal...

Part A

Calculate the pH of 0.85 M NH3 (Kb=1.8×10−5).

Express your answer using two decimal places.

pH=

Part B

Calculate the concentrations of all species present in 0.85 M NH3 (Kb=1.8×10−5).

Express your answers using two significant figures separated by commas. Enter the concentrations of the species in the order listed below.

[NH4+], [NH3], [OH−], [H3O+] =
Answer #1

Part A , pH of the solution is 11.6

Part B , [NH4+] = 3.91 × 10-3 M ; [NH3] = 0.84609 ; [OH-] = 3.91×10-3 M ; [H3O+] = 2.557×10-12 M .